hi i have a question that reads 81+17x=30 and i have an answere of -3 can you tell me how to arrive at this answere as i can only get -2.133 thanx......
Answer by askmemath(368)

81+17x=30
Subtract 81 on both sides
81+17x-81=30-81
17X = -51
Dividing by 17 throughout
X = -51/17 = -3